# TESSERA: Precomputed FAIR Global Pixel Embeddings for Earth Representation and Analysis

### v5

@@ 1 Introduction
Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation [1, 2, 3]. Although petabytes of EO data are available from sensors in various modalities [4], practitioners face fundamental bottlenecks: the scarcity of large, accurately labelled datasets required for supervised training [5], and poor data quality due to sensor noise, cloud cover, atmospheric effects, and non-uniform temporal sampling [6, 7, 8].
To mitigate these issues, users typically employ annual or seasonal compositing [9], then apply supervised machine learning to map from composited images to task-specific classification [10]. However, compositing removes the critical temporal signal arising from phenology, and training task-specific ML models is computationally onerous.
Recent EO approaches address the labelling problem using foundation models: general-purpose models pre-trained on broad EO datasets [11]. These include supervised pre-training on large labelled datasets [12, 13, 14, 15], and self-supervised learning (SSL) strategies that learn from data structure itself. Common SSL techniques include masked autoencoders [16, 17, 18, 19], diffusion-based models [20, 21, 22], and contrastive methods [23, 24, 25, 26]. These foundation models require fine-tuning for specific downstream tasks such as habitat mapping and semantic segmentation [27, 28, 29, 30, 31].
Although existing SSL approaches create rich multidimensional embeddings from unlabelled data [32, 33], they still do not address temporal signal loss during compositing. Lisaius et al. [5] demonstrated that the Barlow Twins loss [34], grounded in redundancy reduction principles [35], can preserve spectral-temporal signals in latent embeddings when processing dense EO time series.
Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10 m resolution and global scale. TESSERA outperforms state-of-the-art methods across diverse tasks including estimating canopy height in Borneo rainforests, detecting crop types in Austria, mapping above-ground biomass in Finland, identifying fire scars in California and evaluating agroforestry stocking indices in Brazil.
Critically, TESSERA provides precomputed outputs that are fully compliant with FAIR (Findable, Accessible, Interoperable, and Reusable) data principles [36]. The embeddings are findable through a per-tile version-controlled registry, openly accessible without requiring inference, interoperable via standard numpy arrays and GeoTIFF metadata, and reusable under an open license with fully reproducible model training from open Sentinel-1/2 data. Users need neither fine-tune the model, download voluminous satellite data, nor perform inferencing beyond downstream tasks. This computational frugality, data privacy preservation, and FAIR compliance removes crucial barriers to use. TESSERA is label-efficient, requiring only small numbers of labels for task-specific training to achieve state-of-the-art performance, and is open source for extensibility and reproducibility.
We provide a gap-free, global, annual, 10 m resolution embedding map for 2024 through the Geotessera Python library [37], with plans for global maps covering 2017-2024. We have also released interactive and batch tools for habitat mapping [38].
TESSERA is unprecedented in combining ease of use, scale, and accuracy: no other foundation model provides precomputed outputs while being open and offering global, annual coverage at 10 m resolution using pixel-level spectral-temporal features (Supplementary Text Table 5). The closest work is Google’s AlphaEarth Satellite Embedding foundation model (GSE) [39], which provides similar global coverage but is neither open nor reproducible. We demonstrate that TESSERA embeddings match or outperform AlphaEarth embeddings in several downstream tasks.
2 Methods
Figure 1: Data processing pipeline. a, Overview. Time-series data from Sentinel-2 MSI and Sentinel-1 SAR observations are extracted for each pixel location to form mode-specific ‘d-pixels’ ( vertical rectangles; cloud-corrupted rows in gray). Two augmentations are created from each d-pixel as cloud-free random samples from the corresponding d-pixel. Augmentations are processed through a dual-branch architecture with weights shared between the branches. The two resulting 128-dimensional embeddings are fused by a multilayer perceptron (MLP) to create the final output embeddings. During training (but not inferencing) the fused features are expanded to 16,384 dimensions by a large projector network. A modified Barlow Twins loss (a linear combination of $\mathcal{L}_{BT}$ and $\mathcal{L}_{MIX}$) is calculated from the cross-correlation matrix of projected features, forcing the model to learn dimensionally decorrelated embeddings that are invariant to masked days in the Sentinel-2 d-pixel. b, Detailed structure of a single encoder branch. Each d-pixel is embedded using a single-layer MLP (not shown) and a Day-of-Year-based temporal encoding is added. The resulting vector is processed by an 4-block Transformer Encoder and then a Gated Recurrent Unit (GRU) pooling layer; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
TESSERA uses self-supervised learning to create per-pixel embeddings from a global dataset of unlabelled time-series of observations from Sentinel-1 Synthetic Aperture Radar (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams. It is trained by optimizing a modified Barlow Twins loss function [34]. The original Barlow Twins approach seeks to create embeddings that are a) invariant to augmentations of its input, where these augmentations are typically rotations, reflections, and cropping of an image and b) whose dimensions are as decorrelated as possible. Inspired by the work by Lisaius et al. [5], we use the Barlow Twins approach to learn embeddings that are invariant to the choice of random subsets of cloud-free observations in the Sentinel-2 optical time series, fused with the corresponding Sentinel-1 data, while minimizing redundancy between the learned feature dimensions. This is achieved by randomly sampling non-cloudy days from the Sentinel-2 annual observation time series so that model then learns to reconcile these partial views, interpolating missing data.
The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Supplementary Text Equation 11). We now discuss the system architecture.
TESSERA comprises two main components: a dual-branch encoder and a projector network (see Fig. 1 for details). TESSERA models its input as a set of ‘d-pixels.’ Each d-pixel represents the annual time series of Sentinel-1 and Sentinel-2 observations at a specific 10 m pixel on Earth as a two-dimensional array ordered in time (that is, channels $\times$ timesteps). This format preserves the temporal evolution of the spectral/backscatter signatures inherent in the satellite data, while allowing observational gaps and irregularities, such as those due to cloud cover or orbit overlap. Rows that are cloud-corrupted or otherwise invalid are indicated by a binary mask and ignored when creating augmentations. Model training and inferencing is done by random temporal sampling followed by flattening of these d-pixels (see Fig. 1 and Supplementary Text for details). In practice, we randomly sub-sample each d-pixel three times, to create a total of three pairs of augmentations. For each 10 m pixel, the fused TESSERA encoders generate a compact (8 bit), 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics. Under typical mid-latitude observation conditions, this compresses the input data by a factor of $\approx$ 24, with higher compression ratios at higher latitudes.
TESSERA uses two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data in the C-band (VV and VH polarizations) and another for Sentinel-2 MSI data (10 spectral bands covering the visible, near, and shortwave infrared). Each encoder processes a d-pixel, utilizing multi-head self-attention mechanisms to capture the temporal dependencies and patterns specific to each data modality. Positional encodings derived from the Day-of-Year (DoY) are incorporated to provide temporal context. An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality [38]. See Figure 1 and Supplementary Text for architectural details.
Modality-specific embeddings derived from dual encoders are fused using a multilayer perceptron (MLP) which is trained alongside the encoders. This fused embedding, which serves as the 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network. This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate redundancy reduction during the self-supervised learning phase. A detailed breakdown of the model’s architectural parameters is provided in the Supplementary table 2.
We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance  [38].
2.1 Model Training
The pre-training phase utilizes a large multi-year global dataset of approximately 800 million d-pixels, down-sampled 400:1 from Sentinel-1 and Sentinel-2 time series from the time period 2017-2024. To ensure successful generalization, a multilevel data sampling, shuffling, and batching strategy was developed. This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features. The model was trained for one epoch in this dataset using distributed training techniques (see [38] for further details on data preparation, the shuffling strategy, and optimization). The key hyperparameters used during pre-training are summarized in Supplementary table 4.
Figure 2: TESSERA embeddings capture fine-grained geospatial patterns across multiple scales. The first three dimensions of the 128-dimensional embeddings are mapped to RGB channels for visualization. a, Examples from diverse global locations, showing a variety of landforms and vegetation patterns. b, A multi-scale view of a complex agricultural landscape. b.1 shows the overall region, with subsequent panels (b.2, b.3) zooming into the white-boxed areas, revealing progressively finer details down to individual crop fields. c, Continental-scale visualizations demonstrating the richness and consistency of the embeddings over vast areas, including (c.1) the Continental United States, (c.2) Europe, and (c.3) the Amazon basin.
2.2 Model Inference and Global Embeddings Maps
Following pre-training, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10 m pixel globally. This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations. The primary output of this inference stage is a set of annual, global, 10 m resolution TESSERA embedding maps. These maps are designed as readily usable multichannel geospatial data layers. This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Supplementary Text). We are currently generating global embeddings for 2024 (approx. 250 TB) and will progressively extend coverage backward year by year until 2017. Visualizing the first three dimensions of the embeddings as R, G, and B channels reveals rich, detailed textures corresponding to diverse land cover and land use patterns. These visualizations highlight the model’s ability to capture information across multiple scales, from individual agricultural fields to entire continents (Fig. 2).
3 Downstream Tasks
The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks. These include pixel-wise classification, such as for crop type mapping, pixel-wise regression, such as for estimating canopy height and carbon density of forests, and patch-based dense predictions, such as for semantic segmentation. (See Supplementary fig. 12 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings. Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these embeddings, acting as explanatory variables in regression modelling. This approach is computation- and labelled-data efficient: it typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state-of-the-art.
3.1 Crop Type Classification
Figure 3: TESSERA embeddings achieve state-of-the-art performance in crop type classification. The evaluation was conducted on the 2022 Austrian crop dataset. a, Weighted and Macro F1 scores for pixel-wise classification as a function of the training data ratio. TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model and other foundation models like PRESTO and Google AlphaEarth Satellite Embedding (GSE). b, Patch-based semantic segmentation performance comparison. TESSERA embeddings, used with a UNet model, achieve higher mean Intersection over Union (mIoU) and Macro F1 scores than other foundation models across various patch sizes and training data ratios. c, Weighted F1 score in few-shot learning scenarios. TESSERA maintains a significant performance advantage even with very few training samples per class. Error bars represent the standard deviation over multiple runs. d, UMAP visualization of the embedding spaces for TESSERA, PRESTO, and GSE for 17 crop classes. TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.
Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [40]. The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions [7]. However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes [7]. In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard [7, 41]. Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally [42, 43]. Here, we demonstrate that TESSERA’s precomputed embeddings set a new state-of-the-art.
To evaluate performance, we used the INVEKOS Austrian crop dataset [44] from the 2021–2022 growing season. This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [44]. We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [19], and (3) Google AlphaEarth Satellite Embedding (GSE) [39]. For the foundation models, we used the same small neural network to classify their respective embeddings to allow a direct comparison. We evaluated performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labelled data to one-shot learning. We also evaluated performance on a patch-based semantic segmentation task against several other foundation models. Details on data processing, model architectures, and training protocols are provided in the Supplementary Text.
TESSERA consistently surpasses both baselines in all training regimes, especially in settings where labelled data are scarce (Fig. 3). In pixel-wise classification tasks with training data splits from 1% to 30%, a simple MLP trained on TESSERA embeddings achieves higher weighted and macro F1 scores than an identical MLP trained on PRESTO or GSE embeddings, and substantially outperforms the RF baseline (Fig. 3a). TESSERA also maintains statistically significant advantages in one-shot and few-shot scenarios (Fig. 3c). Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models, including PRESTO and GSE, across various patch sizes (Fig. 3b).
These performance gains are due to the high quality of the learned embeddings. A UMAP analysis [45] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig. 3d). This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.
3.2 Canopy Height Estimation
Figure 4: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation. Predicted canopy height (y-axis) is compared against airborne LiDAR reference data (x-axis) across multiple models and input datasets in Danum Valley, Borneo. Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google AlphaEarth Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure. TESSERA achieves the highest coefficient of determination and the lowest RMSE ($8.88\pm 0.98$ m), indicating superior performance in structurally complex, tall tropical canopies. Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM. All four products show substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14$ m), with three exhibiting negative correlations. These models saturate at lower canopy heights and show large bias and RMSE, highlighting the limitations of global products when applied to high-biomass tropical forests without local calibration. A 1:1 reference line (blue) is shown in each panel for comparison.
Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+ [46, 47]. It is correlated with ecosystem function [48], forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness [49, 50, 51]. Accurate, scalable canopy height models (CHMs) can enable detection of forest degradation and disturbance.
Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability. Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range [52]. GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes [53], and does not produce unreliable estimates in mountainous terrain [54]. Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [55]. Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [56, 57] - yet none of the three most widely used global CHM models currently implement sensor fusion [58, 59, 60]. Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500 features in Potapov et al. [59]), which may generalize poorly across ecosystems. These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [61], and inconsistencies in input quality. As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes [62]. Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration. These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [63, 64].
We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo [65] with openly available ground-truth data [66]. To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation. A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input. To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set. The accuracy metrics for the foundation models are reported as the average across the 12 runs. TESSERA achieved a coefficient of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62 m.
We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig. 4). First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al. [59], the ETH map by Lang et al. [58], and the Meta map by Tolan et al. [60]. Although local regression models, such as ours, generally outperform global products in regional evaluations [67], the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages [59, 58, 60]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
We also compared our results with regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al. [68], which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass. Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [19] and Google AlphaEarth Satellite Embedding (GSE) [39], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models. Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
In all comparisons, TESSERA outperformed competing methods. Moreover, its embeddings do not require pre-processing, greatly enhancing reproducibility. Instead, only the region and year of interest are needed to replicate the input data pipeline. By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or careful selection of input features, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
3.3 Above-ground Forest Biomass Estimation
Figure 5: A regressor trained on TESSERA embeddings matches the performance of winner of the BioMassters competition with far fewer labels, outperforms SpectralGPT+UPerNet, the best performing foundation model in the BioMassters challenge, and matches the performance of the commercially-developed GSE embeddings. a, TESSERA achieves lower RMSE values than SpectralGPT+UPerNet and results similar to the state-of-the-art supervised baseline and Google Satellite Embedding foundation model (top row). The coefficient of determination ($R^{2}$) and mean bias (MB) are shown in the middle and bottom row as functions of label availability. b, Mean predicted AGB values show increasing bias at higher ground-truth AGB values, which, however, are sparse in the dataset (the y-axis in the upper panel is on a logarithmic scale). Mean predictions are calculated for ground-truth AGB values grouped into 5 t/ha intervals. c, Predicted AGB map for an example patch using 4% of labels in training. d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (left) and the supervised baseline (right). The supervised baseline ignores ground-truth values above 400 t/ha during training, which is reflected in its lower maximum predictions.
Forests play a central role in the global carbon cycle [69]. This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery [70]. A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions. Thus, there has been a long-standing effort to link field measurements to global satellite observations [71].
Recently, with the creation of the BioMassters benchmark dataset [72] (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice  [73]. We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98. In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT [18] and GSE [39].
First, a summary of the BioMassters dataset. This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations [74, 75] informed by extensive field plots. Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig. 5b) that would not appear in coarser resolution data. To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
We compared a UNet-based regression model [76] trained on TESSERA embeddings with SpectralGPT embeddings [18] combined with a lightweight spatial-temporal encoder [77] for temporal aggregation and a UPerNet [78] regression head, because this combination has been found in the Pangaea benchmark [79] to outperform other foundation models [72, 79]. Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase. Thus, we adopted the Pangaea implementation of their model [79], which uses preprocessed satellite imagery from the BioMassters dataset. We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline. This model uses a UNet with a temporal attention encoder [72]. Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable. Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
Our performance metrics are the mean RMSE, the coefficient of determination ($R^{2}$), and the mean bias (MB) in AGB across the test patches as a function of the fraction of labelled data used to train the regressor (Fig. 5a). We make three significant findings. First, although custom models developed using supervised learning are typically difficult to outperform, even with a small fraction of training labels, our regressor nearly matches the performance of the BioMassters competition winner, which uses a supervised learning approach. Second, our regressor trained on 1% of the labelled data outperforms a UPerNet-based regressor trained on the entire training set using SpectralGPT embeddings. Third, our regressor achieves performance comparable to GSE, a commercially developed foundation model with far more training resources at its disposal.
3.4 Burned Area Detection
Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California. a, UMAP projection of $20$ TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events. A continuous gradient within the August fire aligns with MTBS burn severity classes, suggesting that TESSERA captures both timing and magnitude of disturbance. b, Burned vs. unburned classification performance using random forest classifiers trained on different input representations. TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google AlphaEarth Satellite Embeddings (GSE). c, Burn classification maps generated from models trained on 1% of labelled data demonstrate strong agreement with MTBS perimeters across both test areas. As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.
The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [80, 81, 82]. Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [83], predicting wildfire occurrence [84], assessing ecosystem recovery [85], and establishing effective fire management strategies [86].
Remote sensing is widely used to detect burned areas and assess severity regionally and globally [87]. Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [87, 88]. These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [87]. Most existing methods use optical sensors [87], which can detect fire-specific spectral signatures but are susceptible to cloud cover [89], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [90]. SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [87, 91, 92]. Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness [87, 93]. Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [91, 92, 94]. TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig. 6).
We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Fig. 6). Our analysis focused on two randomly selected $\times15$ km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset [95]. UMAP projections [45] of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters. In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Fig. 6a). To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels. We compared three input representations: (1) TESSERA embeddings, (2) Google AlphaEarth Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post- fire Sentinel-2 composites. The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Fig. 6b). Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency. These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps. These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Fig. 6c). The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
3.5 Stocking Indices in Voluntary Carbon Markets
Figure 7: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market (data has been linearly rescaled). Comparison of in-situ AGB measurements in 38 agro-forestry sites, within which 1-3 inventory plots were measured in Para State, Brazil with TESSERA (top left) and five other remote sensing products. Note that the different products were originally provided in different units and converted to AGB (see  [38] for details).
The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere [96]. To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas. This is a costly and inefficient undertaking if it is done only by field sampling.
[…]

### v6

@@ 1 Introduction
Figure 1: TESSERA advances the embedding-as-data approach for Earth Observation. It delivers (a) analysis-ready global-scale products, (b) pixel-wise, high-fidelity representations, and has (c) state-of-the-art downstream accuracy.
Earth Observation (EO) satellites continuously observe our planet, yet cloud cover, irregular revisit intervals depending on latitude, sensors that differ in repeat period, spatial and spectral resolutions, and noise complicate the resulting time series. For many EO applications, such as agriculture, forestry, and environmental monitoring, clouds frequently obscure the evolution of target phenomena. Although compositing can yield cloud-free mosaics and reduce noise in microwave measurements, it flattens phenological dynamics and transient events that downstream models require [114].
Recent remote-sensing foundation models (RSFMs) have made substantial progress in large-scale pretraining, yet most assume heavily preprocessed cloud-filtered inputs. These predominantly patch-based models train on composites or temporal averages, a simplification that discards partly-clouded observations that contain valuable temporal information. Consequently, learned embeddings often bias toward idealized, cloud-free conditions, limiting their robustness to the irregular, incomplete sampling that characterizes operational EO time series [107].
We approach learning EO-based embeddings differently: rather than filtering imperfect observations, we use Barlow Twins [113, 52] to create embeddings that are robust to temporal sampling variability. Specifically, by forcing embeddings to be invariant to randomly selected subsets of cloud-free observations from the same location, our embeddings model consistent physical processes, which results in their generalization across sensors, seasons, and regions.
Building on work by Lisaius et al. [52], which describes the use of Barlow Twins for EO for optical data streams, we introduce TESSERA, a pixel-wise RSFM for multi-modal (Sentinel-1/2) sequences. Our work differs from that of Lisaius et al. [52] in four additional ways. First, their work created embeddings from Sentinel 2 alone, whereas TESSERA combines observations from Sentinel 1 and 2 in a novel architecture. Second, we introduce two complementary regularizers: global shuffling to disrupt short-range spatial correlations, while mix-based regularization blends observations across time and samples to stabilize learning under severe sparsity.
Third, unlike nearly all prior work, open-source global, annual, pixel-wise int8 embeddings for each 10m $\times$ 10m location on land, along with open-source model weights, code, and the GeoTessera Python library***https://github.com/ucam-eo/geotessera to efficiently retrieve embeddings from our data store for large-scale inference. TESSERA embeddings can be paired with lightweight adaptation heads (small MLPs or shallow convolutional models) to achieve high performance with minimal computation, enabling rapid prototyping and straightforward deployment. TESSERA thus emphasizes usability through the FAIR principles (Findable, Accessible, Interoperable, and Reusable).
Finally, we evaluate TESSERA on a wide range of classification, segmentation, and regression benchmarks spanning multiple regions and label regimes. We find that our embeddings remain consistent when calendars shift, seasons are partially missing, or cloud cover is dense. Moreover, we achieve state-of-the-art accuracy with strong label efficiency, particularly under low-label conditions, and the TESSERA embeddings degrade gracefully under heavy cloud cover. Ablations confirm that temporal sampling invariance, global shuffling, and mix-based regulation contribute each to robustness and generalization.
Our contributions are:
  1. 1.
    Global, pixel-wise, label-efficient embeddings. We use a novel self-supervised architecture to train a multi-modal (S1/S2) pixel-wise RSFM.
  2. 2.
    FAIR embeddings-as-data. We are releasing global, annual, 10 m, pixel-wise int8 embeddings with open-source weights, code, and a Python library—advancing FAIR deployment-ready resources.
  3. 3.
    Comprehensive evaluation. We compare our embeddings with a large number of other state-of-the-art foundation models and find that, for diverse classification, segmentation, and regression tasks, TESSERA embeddings deliver outstanding performance.
2 Related Work
2.1 Foundation Models for Earth Observation
Remote-sensing foundation models (RSFMs) have rapidly advanced along several methodological lines. A large body of work has used self-supervised pretraining using either contrastive objectives (e.g., SeCo [59], SkySense [31]) or masked image modeling (e.g., SatMAE [15], Scale-MAE [71], Cross-Scale MAE [84], S2MAE [47], SatMAE++ [65]). Other variants leverage global–local view alignment or distilled objectives (e.g., DINO-MM/DINO-MC [98, 105]). Model scopes have expanded from single-modality encoders toward broader families such as RingMo and RingMo-Lite [81, 100], Satlas [8], Prithvi and Prithvi-EO-2.0 [76, 82], GFM/msGFM [61, 32], Skysense series [31, 106, 56, 121, 119]. Work on multi-sensor fusion explicitly integrates optical and SAR (e.g., CROMA [29], SkySense [31], RingMo-Sense [112]). In parallel, early efforts like MOSAIKS [72] explored lightweight universal features to lower adoption barriers.
Despite these advances, most RSFMs are trained on patch-based inputs and implicitly assume that the data have been preprocessed (e.g., aggressive cloud filtering or temporal averaging). Such assumptions simplify pretraining but discard partially cloudy images that still carry useful temporal signals, biasing representations toward idealized, cloud-free conditions. Moreover, strong performance typically requires fine-tuning the backbone for each downstream task, which can be compute- and label-intensive for many EO users.
2.2 Embedding-Based Foundation Models
A complementary line of work delivers embeddings-as-data: precomputed, generic, task-agnostic features that can be easily adapted with lightweight heads (shown in Figure 2 c). Representative models include Presto [88], which learns pixel-level temporal representations, and Google Deepmind’s AlphaEarth [11], which provides precomputed embeddings as a ready-for-analysis product. This paradigm reduces computational costs and democratizes access to high performance; however, existing approaches are often not fully open-sourced, lack pixel-level granularity at a global scale, or do not explicitly account for irregular temporal sampling.
[…]